OpenStudy (anonymous):

what is the surface area of a sphere with a volume of 140

OpenStudy (amistre64):

Sa = 4pi r^2 V = (4pi/3) r^3

OpenStudy (amistre64):

r = cbrt(3V/4pi) ..... maybe ☺

OpenStudy (anonymous):

options are 880 130 420 339

OpenStudy (anonymous):

130!

myininaya (myininaya):

Volume of sphere is 4/3 *pi*r^3=140 Surface area of a sphere is 4*pi*r^2=3/pi*(4/3*pi*r^3)=3/pi*(140)

OpenStudy (amistre64):

130 if i did it right

myininaya (myininaya):

i got 133.69 blah blah

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

thanks yall!

OpenStudy (anonymous):

did u know the answer??????

OpenStudy (anonymous):

:~\

OpenStudy (anonymous):

4/3pir^2